# homogeneous_khovanskii.bib

### v1

@@ 2.3. Computing homogeneous Khovanskii bases
Let $f_{1},\dots,f_{k}\in A$ be $\Delta$/̄homogeneous generators of $A$ as a $K$/̄algebra. We consider the two morphisms
@@
We endow the polynomial ring $K[X_{1},\dots,X_{k}]$ with a grading by $\Delta$ via $\deg_{\Delta}(X_{i})\coloneqq\deg_{\Delta}(f_{i})$ and a grading by $\Gamma$ via $\deg_{\Gamma}(X_{i})\coloneqq v(f_{i})=\deg_{v}(\operatorname{in}_{v}^{\Delta}(f_{i}))$. The morphisms $\alpha$ and $\beta$ are $\Delta$/̄graded by construction and $\beta$ is a $\Gamma$/̄graded morphism.
Notation 2.14.
We write $\operatorname{in}_{\Gamma}(h)\in K[X_{1},\dots,X_{k}]$ for the sum of the terms of minimal $\Gamma$/̄degree of a polynomial $h\in K[X_{1},\dots,X_{k}]$ (with respect to $\succeq$).
@@
Given a $\Delta$/̄homogeneous polynomial $h\in I$, the initial form $\operatorname{in}_{\Gamma}(h)$ is $\Delta$/̄homogeneous as well. Hence $\operatorname{in}_{\Gamma}(I)$ is a $\Delta$/̄homogeneous ideal because $I$ is $\Delta$/̄homogeneous. As also $J$ is $\Delta$/̄homogeneous, it suffices to check the claim on $\Delta$/̄homogeneous elements.
The remainder of the proof is now analogous to [KM19, Lemma 2.16]. Let $h\in I$ be $\Delta$/̄homogeneous and set $\gamma\coloneqq\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h))$. Then $v(\operatorname{in}_{\Gamma}(h)(f_{1},\dots,f_{k}))\succ\gamma$ as $v(\operatorname{in}_{\Gamma}(h)(f_{1},\dots,f_{k}))=\gamma$ would contradict $h(f_{1},\dots,f_{k})=0$. Hence
@@
by Lemma 2.15. By $\Delta$/̄homogeneity, we conclude $\beta(\operatorname{in}_{\Gamma}(h))=0$, so $\operatorname{in}_{\Gamma}(h)\in J$. ∎
@@
Let $0\neq h\in K[X_{1},\dots,X_{k}]$ be $\Gamma$- and $\Delta$/̄homogeneous. We have $h\in\operatorname{in}_{\Gamma}(I)$ if and only if there exists a $\Delta$/̄homogeneous polynomial $\tilde{h}\in K[X_{1},\dots,X_{k}]$ such that $h-\tilde{h}\in I$ and $\deg_{\Gamma}(h)\prec\deg_{\Gamma}(\operatorname{in}_{\Gamma}(\tilde{h}))$.
@@
Assume $h\in\operatorname{in}_{\Gamma}(I)$, so we can write $h=\sum_{i=1}^{t}\operatorname{in}_{\Gamma}(h_{i})$ with $h_{i}\in I$. We may assume that the $h_{i}$ are $\Delta$/̄homogeneous of same degree since $h$ is $\Delta$/̄homogeneous and $I$ is a $\Delta$/̄homogeneous ideal. Set $h^{\prime}\coloneqq\sum_{i=1}^{t}h_{i}\in I$. Then $\operatorname{in}_{\Gamma}(h^{\prime})=h$ by $\deg_{\Gamma}$/̄homogeneity of $h$. Hence $\tilde{h}\coloneqq h-h^{\prime}$ fulfils the requirements.
@@
Let $h\in J$ be $\Gamma$- and $\Delta$/̄homogeneous. If Algorithm 1 terminates with remainder 0 for $h(f_{1},\dots,f_{k})$ and $\{f_{1},\dots,f_{k}\}$, then $h\in\operatorname{in}_{\Gamma}(I)$.
@@
Set $f\coloneqq h(f_{1},\dots,f_{k})$. We have $h(\operatorname{in}_{v}(f_{1}),\dots,\operatorname{in}_{v}(f_{k}))=0$, so $\deg_{\Gamma}(h)\prec v(f)$. By assumption, Algorithm 1 produces an element $h^{\prime}\in K[X_{1},\dots,X_{k}]$ with $f=h^{\prime}(f_{1},\dots,f_{k})$. Further, by (b) in Algorithm 1, we have $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h^{\prime}))\succeq v(f)$, from which we conclude $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h^{\prime}))=v(f)$. Hence $h-h^{\prime}\in I$ and $\operatorname{in}_{\Gamma}(h-h^{\prime})=\operatorname{in}_{\Gamma}(h)=h$, so $h\in\operatorname{in}_{\Gamma}(I)$. ∎
@@
Assume that Algorithm 1 terminates for $A$. Let $f_{1},\dots,f_{k}\in A$ be $\Delta$/̄homogeneous generators. The following conditions are equivalent:
@@
    The morphism $\beta$ is surjective, that is, $\{f_{1},\dots,f_{k}\}$ is a $\Delta$/̄homogeneous Khovanskii basis of $A$.
@@
    For a set of generators $\{h_{1},\dots,h_{s}\}$ of $J$ consisting of $\Gamma$/̄homogeneous as well as $\Delta$/̄homogeneous polynomials, Algorithm 1 terminates with remainder 0 for $h_{i}(f_{1},\dots,f_{k})$ and $\{f_{1},\dots,f_{k}\}$, $1\leq i\leq s$.
@@
‘(i)$\Longrightarrow$(ii)’: Let $\beta$ be surjective. Then Algorithm 1 terminates with remainder 0 for every $f\in A$. So, $\operatorname{in}_{\Gamma}(I)=J$ by Lemma 2.16 and Lemma 2.18.
‘(ii)$\Longrightarrow$(i)’: Assume $\operatorname{in}_{\Gamma}(I)=J$ and let $0\neq f\in A$ be $\Delta$/̄homogeneous. We have to show that there is $h\in K[X_{1},\dots,X_{k}]$ with $\beta(h)=\operatorname{in}_{v}^{\Delta}(f)$. Let $h^{\prime}\in K[X_{1},\dots,X_{k}]$ be $\Delta$/̄homogeneous with $h^{\prime}(f_{1},\dots,f_{k})=f$. If $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h^{\prime}))=v(f)$, then $\operatorname{in}_{\Gamma}(h^{\prime})(\operatorname{in}_{v}(f_{1}),\dots,\operatorname{in}_{v}(f_{k}))=\operatorname{in}_{v}(f)$ by Lemma 2.15, so $\operatorname{in}_{\Gamma}(h^{\prime})$ is the desired preimage of $\operatorname{in}_{v}^{\Delta}(f)$ under $\beta$. Assume hence $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h^{\prime}))\prec v(f)$, so $\operatorname{in}_{\Gamma}(h^{\prime})(\operatorname{in}_{v}(f_{1}),\dots,\operatorname{in}_{v}(f_{k}))=0$ by Lemma 2.15 and $\operatorname{in}_{\Gamma}(h^{\prime})\in J$. By assumption, $\operatorname{in}_{\Gamma}(h^{\prime})\in\operatorname{in}_{\Gamma}(I)$, so there is $\tilde{h}\in K[X_{1},\dots,X_{k}]$ with $\operatorname{in}_{\Gamma}(h^{\prime})-\tilde{h}\in I$ and $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h))\prec\deg_{\Gamma}(\operatorname{in}_{\Gamma}(\tilde{h}))$ by Lemma 2.17. Hence for $h^{\prime\prime}\coloneqq h^{\prime}-\operatorname{in}_{\Gamma}(h^{\prime})+\tilde{h}$ we have $\alpha(h^{\prime\prime})=f$ and $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h^{\prime\prime}))\succ\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h^{\prime}))$. Increasing the $\Gamma$/̄degree of the initial form of a preimage of $f$ in this way, we eventually obtain a preimage $h\in\alpha^{-1}(f)$ with $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h))=v(f)$. Then $\operatorname{in}_{\Gamma}(h)$ gives the desired preimage under $\beta$ as above.
‘(i)$\Longrightarrow$(iii)’: By assumption, Algorithm 1 terminates for $A$. As $\beta$ is surjective, the algorithm must return remainder 0 for every $f\in A$.
‘(iii)$\Longrightarrow$(ii)’: Let $J=\langle h_{1},\dots,h_{s}\rangle$ with $\Delta$/̄homogeneous and $\Gamma$/̄homogeneous polynomials $h_{i}$ and assume Algorithm 1 reduces $h_{i}(f_{1},\dots,f_{k})$ to 0. This implies $h_{i}\in\operatorname{in}_{\Gamma}(I)$ by Lemma 2.18. Hence $J\subseteq\operatorname{in}_{\Gamma}(I)$ and the claim follows with Lemma 2.16. ∎
@@
Input : $\Delta$/̄homogeneous generators $f_{1},\dots,f_{k}\in A\leq B$
Output : A $\Delta$/̄homogeneous Khovanskii basis of $A$ with respect to $v$
1 Initialize $\mathcal{B}\coloneqq\mathcal{B}^{\prime}\coloneqq\{f_{1},\dots,f_{k}\}$
2 while $\mathcal{B}^{\prime}\neq\emptyset$ do
3      $\mathcal{B}^{\prime}\coloneqq\emptyset$
4      With =$\mathcal{B}=\{f^{\prime}_{1},\dots,f^{\prime}_{l}\}$, set up the morphism
@@
and compute $\Gamma$/̄homogeneous as well as $\Delta$/̄homogeneous generators of the kernel $\ker(\beta)=\langle h_{1},\dots,h_{s}\rangle$
@@
6           if Algorithm 1 terminates for $h(f_{1}^{\prime},\dots,f^{\prime}_{l})$ and $\mathcal{B}$ with a non-zero remainder $r$ then
7                $\mathcal{B}^{\prime}\coloneqq\mathcal{B}^{\prime}\cup\{r\}$
8
@@ 7                $\mathcal{B}^{\prime}\coloneqq\mathcal{B}^{\prime}\cup\{r\}$
12               $\mathcal{B}\coloneqq\mathcal{B}\cup\mathcal{B}^{\prime}$
13
@@ 12               $\mathcal{B}\coloneqq\mathcal{B}\cup\mathcal{B}^{\prime}$
Assume that Algorithm 1 terminates for $A$. Algorithm 2 terminates after finitely many steps if and only if there exists a finite $\Delta$/̄homogeneous Khovanskii basis of $A$ with respect to $v$. If the algorithm terminates, it correctly returns such a basis.
@@
Assume that Algorithm 2 terminates with output $\{f_{1}^{\prime},\dots,f_{l}^{\prime}\}$. By construction, these polynomials are generators of $A$ as the set contains the original generators. Further, the $f^{\prime}_{i}$ are $\Delta$/̄homogeneous as Algorithm 1 maintains the homogeneity. Let $\beta:K[X_{1},\dots,X_{l}]\to\operatorname{in}_{v}^{\Delta}(A)$ be the map corresponding to the $f_{i}^{\prime}$ and set $J\coloneqq\ker(\beta)$. As no generators were added in the final iteration of the while/̄loop, there are generators $h_{1},\dots,h_{s}$ of $J$ such that Algorithm 1 returns remainder 0 for $h_{i}(f_{1}^{\prime},\dots,f_{l}^{\prime})$ and $\{f_{1}^{\prime},\dots,f_{l}^{\prime}\}$. Hence $\{f_{1}^{\prime},\dots,f_{l}^{\prime}\}$ is a $\Delta$/̄homogeneous Khovanskii basis of $A$ by Theorem 2.19. In particular, $A$ admits such a finite basis.
Conversely, assume that $A$ contains a finite $\Delta$/̄homogeneous Khovanskii basis $\{\tilde{f}_{1},\dots,\tilde{f}_{s}\}$. Let $f_{1}^{\prime},\dots,f_{l}^{\prime}\in A$ be the polynomials in a given iteration of the algorithm and let $\beta$ be the corresponding map. We assume at first that for all $i\in\{1,\dots,s\}$ we have $\operatorname{in}_{v}^{\Delta}(\tilde{f}_{i})\in\operatorname{im}(\beta)$. Then $\beta$ must be surjective as $\operatorname{in}_{v}^{\Delta}(\tilde{f}_{i})$ generate $\operatorname{in}_{v}^{\Delta}(A)$ by construction. Hence $\{f^{\prime}_{1},\dots,f^{\prime}_{l}\}$ is a $\Delta$/̄homogeneous Khovanskii basis as well. By Theorem 2.19, the generators of $\ker(\beta)$ chosen by the algorithm must reduce to 0 in Algorithm 1 and the algorithm terminates.
Assume now that there is some $f\in\{\tilde{f}_{1},\dots,\tilde{f}_{s}\}$ with $\operatorname{in}_{v}^{\Delta}(f)\notin\operatorname{im}(\beta)$. As $f\in A$, there is a $\Delta$/̄homogeneous polynomial $h\in K[X_{1},\dots,X_{l}]$ with $h(f_{1}^{\prime},\dots,f_{l}^{\prime})=f$. By assumption, $\operatorname{in}_{\Gamma}(h)(\operatorname{in}_{v}(f_{1}^{\prime}),\dots,\operatorname{in}_{v}(f_{l}^{\prime}))\neq\operatorname{in}_{v}(f)$, so $\operatorname{in}_{\Gamma}(h)(\operatorname{in}_{v}(f_{1}^{\prime}),\dots,\operatorname{in}_{v}(f_{l}^{\prime}))=0$ by Lemma 2.15. Hence $\operatorname{in}_{\Gamma}(h)\in J$. Let $h_{1},\dots,h_{s}\in K[X_{1},\dots,X_{k}]$ with $J=\langle h_{1},\dots,h_{s}\rangle$ be the generators chosen by the algorithm. For every $i\in\{1,\dots,s\}$, the algorithm computes $h^{\prime}_{i}\in K[X_{1},\dots,X_{l}]$ and $r_{i}\in A$ with
@@
Hence $\operatorname{in}_{\Gamma}(h_{i}-h_{i}^{+})=\operatorname{in}_{\Gamma}(h_{i})=h_{i}$ by (b) in Algorithm 1 and the $\Gamma$/̄homogeneity of $h_{i}$. This implies $h_{i}\in\operatorname{in}_{\Gamma}(I)$. We conclude $\operatorname{in}_{\Gamma}(h)\in\operatorname{in}_{\Gamma}(I)$ and by Lemma 2.17 there is $\tilde{h}\in K[X_{1},\dots,X_{l+s}]$ with $\operatorname{in}_{\Gamma}(h)-\tilde{h}\in I$ and $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h))\prec\deg_{\Gamma}(\operatorname{in}_{\Gamma}(\tilde{h}))$. Then for $h^{+}\coloneqq h-\operatorname{in}_{\Gamma}(h)+\tilde{h}$ we have $h^{+}(f_{1}^{\prime},\dots,f_{l}^{\prime})=f$ and $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h))\prec\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h^{+}))$. Increasing the $\Gamma$/̄degree of a preimage of $f$ in this way, we find a preimage $h$ of degree $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h))=v(f)$ after finitely many iterations. Equivalently, we achieve $\operatorname{in}_{\Gamma}^{\Delta}(f)\in\operatorname{im}(\beta)$ (with an extended map $\beta$) after finitely many iterations. ∎

### v2

@@ 2.3. Computing homogeneous Khovanskii bases
Let $f_{1},\dots,f_{k}\in A$ be $\Delta$-homogeneous generators of $A$ as a $K$-algebra. We consider the two morphisms
@@
We endow the polynomial ring $K[X_{1},\dots,X_{k}]$ with a grading by $\Delta$ via $\deg_{\Delta}(X_{i})\coloneqq\deg_{\Delta}(f_{i})$ and a grading by $\Gamma$ via $\deg_{\Gamma}(X_{i})\coloneqq v(f_{i})=\deg_{v}(\operatorname{in}_{v}^{\Delta}(f_{i}))$. The morphisms $\alpha$ and $\beta$ are $\Delta$-graded by construction and $\beta$ is a $\Gamma$-graded morphism.
Notation 2.16.
We write $\operatorname{in}_{\Gamma}(h)\in K[X_{1},\dots,X_{k}]$ for the sum of the terms of minimal $\Gamma$-degree of a polynomial $h\in K[X_{1},\dots,X_{k}]$ (with respect to $\succeq$).
@@
Given a $\Delta$-homogeneous polynomial $h\in I$, the initial form $\operatorname{in}_{\Gamma}(h)$ is $\Delta$-homogeneous as well. Hence $\operatorname{in}_{\Gamma}(I)$ is a $\Delta$-homogeneous ideal because $I$ is $\Delta$-homogeneous. As also $J$ is $\Delta$-homogeneous, it suffices to check the claim on $\Delta$-homogeneous elements. The remainder of the proof is now analogous to [KM19, Lemma 2.16]. Let $h\in I$ be $\Delta$-homogeneous and set $\gamma\coloneqq\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h))$. We have $h(f_{1},\dots,f_{k})=0$ and hence
@@
by Lemma 2.17. By $\Delta$-homogeneity, we conclude $\beta(\operatorname{in}_{\Gamma}(h))=0$, so $\operatorname{in}_{\Gamma}(h)\in J$. ∎
@@
Let $0\neq h\in K[X_{1},\dots,X_{k}]$ be $\Gamma$- and $\Delta$-homogeneous. We have $h\in\operatorname{in}_{\Gamma}(I)$ if and only if there exists a $\Delta$-homogeneous polynomial $\tilde{h}\in K[X_{1},\dots,X_{k}]$ such that $h-\tilde{h}\in I$ and $)\deg_{\Gamma}(h)\prec\deg_{\Gamma}(\operatorname{in}_{\Gamma}(\tilde{h}$)).
@@
Assume $h\in\operatorname{in}_{\Gamma}(I)$, so we can write $h=\sum_{i=1}^{t}\operatorname{in}_{\Gamma}(h_{i})$ with $h_{i}\in I$. We may assume that the $h_{i}$ are $\Delta$-homogeneous of same degree since $h$ is $\Delta$-homogeneous and $I$ is a $\Delta$-homogeneous ideal. Set $h^{\prime}\coloneqq\sum_{i=1}^{t}h_{i}\in I$. Then $\operatorname{in}_{\Gamma}(h^{\prime})=h$ by $\deg_{\Gamma}$-homogeneity of $h$. Hence $\tilde{h}\coloneqq h-h^{\prime}$ fulfils the requirements.
@@
Let $h\in J$ be $\Gamma$- and $\Delta$-homogeneous. If Algorithm 1 terminates with remainder 0 for $h(f_{1},\dots,f_{k})$ and $\{f_{1},\dots,f_{k}\}$, then $h\in\operatorname{in}_{\Gamma}(I)$.
@@
Set $f\coloneqq h(f_{1},\dots,f_{k})$. We have $h(\operatorname{in}_{v}(f_{1}),\dots,\operatorname{in}_{v}(f_{k}))=0$, so $\deg_{\Gamma}(h)\prec v(f)$. By assumption, Algorithm 1 produces an element $h^{\prime}\in K[X_{1},\dots,X_{k}]$ with $f=h^{\prime}(f_{1},\dots,f_{k})$. Further, by (b) in Algorithm 1, we have $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h^{\prime}))\succeq v(f)$, from which we conclude $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h^{\prime}))=v(f)$ by Lemma 2.17. Hence $h-h^{\prime}\in I$ and $\operatorname{in}_{\Gamma}(h-h^{\prime})=\operatorname{in}_{\Gamma}(h)=h$, so $h\in\operatorname{in}_{\Gamma}(I)$. ∎
@@
Assume that Algorithm 1 terminates for $A$. Let $f_{1},\dots,f_{k}\in A$ be $\Delta$-homogeneous generators. The following conditions are equivalent:
@@
    The morphism $\beta$ is surjective, that is, $\{f_{1},\dots,f_{k}\}$ is a $\Delta$-homogeneous Khovanskii basis of $A$.
@@
    For a set of generators $\{h_{1},\dots,h_{s}\}$ of $J$ consisting of $\Gamma$-homogeneous as well as $\Delta$-homogeneous polynomials, Algorithm 1 terminates with remainder 0 for $h_{i}(f_{1},\dots,f_{k})$ and $\{f_{1},\dots,f_{k}\}$, $1\leq i\leq s$.
@@
[…]
@@
Input : $\Delta$-homogeneous generators $f_{1},\dots,f_{k}\in A\leq B$
Output : A $\Delta$-homogeneous Khovanskii basis of $A$ with respect to $v$
1
21exInitialize $\mathcal{G}\coloneqq\mathcal{G}^{\prime}\coloneqq\{f_{1},\dots,f_{k}\}$
3 while $\mathcal{G}^{\prime}\neq\emptyset$ do
4   $\mathcal{G}^{\prime}\coloneqq\emptyset$
5   With $\mathcal{G}=\{f^{\prime}_{1},\dots,f^{\prime}_{l}\}$, set up the morphism
@@
and compute $\Gamma$-homogeneous as well as $\Delta$-homogeneous generators of the kernel $\ker(\beta)=\langle h_{1},\dots,h_{s}\rangle$
@@
7     if Algorithm 1 terminates for $h(f_{1}^{\prime},\dots,f^{\prime}_{l})$ and $\mathcal{G}$ with a non-zero remainder $r$ then
8       $\mathcal{G}^{\prime}\coloneqq\mathcal{G}^{\prime}\cup\{r\}$
9
@@ 8       $\mathcal{G}^{\prime}\coloneqq\mathcal{G}^{\prime}\cup\{r\}$
13      $\mathcal{G}\coloneqq\mathcal{G}\cup\mathcal{G}^{\prime}$
14
@@ 13      $\mathcal{G}\coloneqq\mathcal{G}\cup\mathcal{G}^{\prime}$
Assume that Algorithm 1 terminates for $A$. Algorithm 2 terminates after finitely many steps if and only if there exists a finite $\Delta$-homogeneous Khovanskii basis of $A$ with respect to $v$. If the algorithm terminates, it correctly returns such a basis.
@@
Assume that Algorithm 2 terminates with output $\{f_{1}^{\prime},\dots,f_{l}^{\prime}\}$. By construction, these polynomials are generators of $A$ as the set contains the original generators. Further, the $f^{\prime}_{i}$ are $\Delta$-homogeneous as Algorithm 1 maintains the homogeneity. Let $\beta:K[X_{1},\dots,X_{l}]\to\operatorname{in}_{v}^{\Delta}(A)$ be the map corresponding to the $f_{i}^{\prime}$ and set $J\coloneqq\ker(\beta)$. As no generators were added in the final iteration of the while-loop, there are generators $h_{1},\dots,h_{s}$ of $J$ such that Algorithm 1 returns remainder 0 for $h_{i}(f_{1}^{\prime},\dots,f_{l}^{\prime})$ and $\{f_{1}^{\prime},\dots,f_{l}^{\prime}\}$. Hence $\{f_{1}^{\prime},\dots,f_{l}^{\prime}\}$ is a $\Delta$-homogeneous Khovanskii basis of $A$ by Theorem 2.21. In particular, $A$ admits such a finite basis.
Conversely, assume that $A$ contains a finite $\Delta$-homogeneous Khovanskii basis $\{\tilde{f}_{1},\dots,\tilde{f}_{s}\}$. Let $f_{1}^{\prime},\dots,f_{l}^{\prime}\in A$ be the polynomials in a given iteration of the algorithm and let $\beta$ be the corresponding map. We assume at first that for all $i\in\{1,\dots,s\}$ we have $\operatorname{in}_{v}^{\Delta}(\tilde{f}_{i})\in\operatorname{im}(\beta)$. Then $\beta$ must be surjective as $\operatorname{in}_{v}^{\Delta}(\tilde{f}_{i})$ generate $\operatorname{in}_{v}^{\Delta}(A)$ by construction. Hence $\{f^{\prime}_{1},\dots,f^{\prime}_{l}\}$ is a $\Delta$-homogeneous Khovanskii basis as well. By Theorem 2.21, the generators of $\ker(\beta)$ chosen by the algorithm must reduce to 0 in Algorithm 1 and the algorithm terminates.
Assume now that there is some $f\in\{\tilde{f}_{1},\dots,\tilde{f}_{s}\}$ with $\operatorname{in}_{v}^{\Delta}(f)\notin\operatorname{im}(\beta)$. As $f\in A$, there is a $\Delta$-homogeneous polynomial $h\in K[X_{1},\dots,X_{l}]$ with $h(f_{1}^{\prime},\dots,f_{l}^{\prime})=f$. By assumption, $\operatorname{in}_{\Gamma}(h)(\operatorname{in}_{v}(f_{1}^{\prime}),\dots,\operatorname{in}_{v}(f_{l}^{\prime}))\neq\operatorname{in}_{v}(f)$, so $\operatorname{in}_{\Gamma}(h)(\operatorname{in}_{v}(f_{1}^{\prime}),\dots,\operatorname{in}_{v}(f_{l}^{\prime}))=0$ by Lemma 2.17. Hence $\operatorname{in}_{\Gamma}(h)\in J$. Let $h_{1},\dots,h_{s}\in K[X_{1},\dots,X_{k}]$ with $J=\langle h_{1},\dots,h_{s}\rangle$ be the generators chosen by the algorithm. For every $i\in\{1,\dots,s\}$, the algorithm computes $h^{\prime}_{i}\in K[X_{1},\dots,X_{l}]$ and $r_{i}\in A$ with
@@
Hence $\operatorname{in}_{\Gamma}(h_{i}-h_{i}^{+})=\operatorname{in}_{\Gamma}(h_{i})=h_{i}$ by (b) in Algorithm 1 and the $\Gamma$-homogeneity of $h_{i}$. This implies $h_{i}\in\operatorname{in}_{\Gamma}(I)$. Recall that $\operatorname{in}_{\Gamma}(h)\in\langle h_{1},\dots,h_{s}\rangle$, so we conclude $\operatorname{in}_{\Gamma}(h)\in\operatorname{in}_{\Gamma}(I)$. By Lemma 2.19, there is $\tilde{h}\in K[X_{1},\dots,X_{l+s}]$ with $\operatorname{in}_{\Gamma}(h)-\tilde{h}\in I$ and $)\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h))\prec\deg_{\Gamma}(\operatorname{in}_{\Gamma}(\tilde{h}$)). Then for $h^{+}\coloneqq h-\operatorname{in}_{\Gamma}(h)+\tilde{h}$ we have $h^{+}(f_{1}^{\prime},\dots,f_{l}^{\prime})=f$ and $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h))\prec\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h^{+}))$. Increasing the $\Gamma$-degree of a preimage of $f$ in this way, we find a preimage $h$ of degree $\deg_{\Gamma}(\operatorname{in}_{\Gamma}(h))=v(f)$ after finitely many iterations. Equivalently, we achieve $\operatorname{in}_{\Gamma}^{\Delta}(f)\in\operatorname{im}(\beta)$ (with an extended map $\beta$) after finitely many iterations. ∎
Example 2.23.
This example is part of the series also containing 2.5 and 2.12. Let $\alpha$ and $\beta$ be the maps corresponding to $\{f_{1},f_{2}\}$, that is, $\alpha(X_{i})=f_{i}$ and $\beta(X_{i})=\operatorname{in}_{v}^{\Delta}(f_{i})$. First of all, we have $I=\ker(\alpha)=\{0\}$, so $\operatorname{in}_{\Gamma}(I)=\{0\}$, but $J=\ker(\beta)=\langle X_{1}^{2}-X_{2}^{2}\rangle$. By Theorem 2.21, this confirms our observation that $\{f_{1},f_{2}\}$ is not a $\Delta$-homogeneous Khovanskii basis.
We run Algorithm 2 with $\mathcal{G}=\{f_{1},f_{2}\}$. We have $J=\langle X_{1}^{2}-X_{2}^{2}\rangle$, so we must run Algorithm 1 with $f_{3}\coloneqq f_{1}^{2}-f_{2}^{2}=4xy$ and $\{f_{1},f_{2}\}$. Because $\operatorname{in}_{v}^{\Delta}(f_{3})=4xy\otimes\bar{0}$, the algorithm immediately terminates with remainder $f_{3}$.
So we set $\mathcal{G}=\{f_{1},f_{2},f_{3}\}$. Computing the relations of the initial forms $\operatorname{in}_{v}^{\Delta}(f_{i})$ yields the same ideal $J$ as before, so Algorithm 2 terminates with the set of generators $\mathcal{G}$.
We confirm the result using Theorem 2.21. For the relations of the polynomials $f_{i}$, we obtain the ideal $I=\langle X_{1}^{2}-X_{2}^{2}-X_{3}\rangle$. The grading by $\Gamma=\mathbb{Z}$ on $K[X_{1},X_{2},X_{3}]$ is given by $\deg_{\Gamma}(X_{i})=v(f_{i})$, so $\deg_{\Gamma}(X_{1})=0$, $\deg_{\Gamma}(X_{2})=0$ and $\deg_{\Gamma}(X_{3})=1$. Hence we have $\operatorname{in}_{\Gamma}(I)=\langle X_{1}^{2}-X_{2}^{2}\rangle$ and this coincides with $J$ as required.